Riverbed Company received the following selected information from its pension plan trustee concerning the operation of the company’s defined benefit pension plan for the year ended December 31, 2017.

January 1, 2017

December 31, 2017

Projected benefit obligation $1,523,000 $1,550,000
Market-related and fair value of plan assets 794,000 1,128,400
Accumulated benefit obligation 1,597,000 1,716,400
Accumulated OCI (G/L)—Net gain 0 (202,300 )

The service cost component of pension expense for employee services rendered in the current year amounted to $77,000 and the amortization of prior service cost was $119,400. The company’s actual funding (contributions) of the plan in 2017 amounted to $255,000. The expected return on plan assets and the actual rate were both 10%; the interest/discount (settlement) rate was 10%. Accumulated other comprehensive income (PSC) had a balance of $1,194,000 on January 1, 2017. Assume no benefits paid in 2017.
Determine the amounts of the components of pension expense that should be recognized by the company in 2017. (Enter amounts that reduce pension expense with either a negative sign preceding the number e.g. -45 or parenthesis e.g. (45).)

Components of Pension Expense

Actual Return on Plan AssetsAmortization of Prior Service CostBenefits PaidContributions to PlanExpected Return on Plan AssetsInterest on Projected Benefit ObligationService Cost

$

Actual Return on Plan AssetsAmortization of Prior Service CostBenefits PaidContributions to PlanExpected Return on Plan AssetsInterest on Projected Benefit ObligationService Cost

Actual Return on Plan AssetsAmortization of Prior Service CostBenefits PaidContributions to PlanExpected Return on Plan AssetsInterest on Projected Benefit ObligationService Cost

Actual Return on Plan AssetsAmortization of Prior Service CostBenefits PaidContributions to PlanExpected Return on Plan AssetsInterest on Projected Benefit ObligationService Cost

SHOW LIST OF ACCOUNTS

Prepare the journal entry to record pension expense and the employer’s contribution to the pension plan in 2017. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

(To record pension expense and employer’s contribution)
Answer #1

Answer (a):

Interest on Projected Benefit Obligation = 1523000 * 10% = $152,300

Expected Return on Plan Assets = 794000 * 10% = $79,400

Components of Pension Expense Service Cost Interest on Projected Benefit Obligation Expected Return on Plan Assets Amortizati

Answer (b):

Increase in PBO = 1550000 - 1523000 = $27,000

Increase in assets = 1128400 - 794000 = 334,400

Pension asset = 334400 - 27000 = $307,400

Journal Entry:

Credit Debit $269,300 $307,400 Account Titles and Explanation Pension Expense Pension Assets Cash Other Comprehensive Income